OpenStudy (anonymous):

Maria invested $5500 in two mutual funds.One fund earned an 8% profit,while the other earned a 1.5% profit.Between the two accounts,Maria made a profit of $180. How much was invested in each account?

OpenStudy (nurali):

let x=amount invested at 8% and (5500-x) amount invested at 1.5% 180=8%x+1.5%(5500-x) 180=0.08x+0.015(5500-x) 180=0.08x+82.5-0.015x 180=0.065x+82.5 97.5=0.065x x=97.5/0.065 x=$1500 invested at 8% $5500-$1500=$4000 invested at 1.5%